# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v7

@@ Appendix E Robustness Analysis: Breakdown Point of the Partial Trimmed Mean
$\frac{1}{n}\max\left(\left\lfloor m+\frac{1}{2}-\frac{K(\delta)}{\sqrt{t-T}}\right\rfloor,0\right)\leq\varepsilon^{*}_{i}(t)\leq\frac{m}{n}\enspace,$
@@
Following the proof in Boyd et al. [2006], by repeatedly conditioning, we obtain
@@ Appendix I Large-scale Experiments
Designing faster gossip algorithms is an interesting direction, and we have also been exploring this aspect. One promising approach involves optimizing the edge sampling strategy based on the graph connectivity constant $c$ (see Xiao and Boyd [2004], Boyd et al. [2004]). While this strategy improves gossip performance in standard averaging tasks, our empirical experiments suggest that its impact on the GoRank algorithm is more limited. This is likely due to the relatively small variation in $c$, which is insufficient to significantly affect a rate in 1/ct1/ct—though it has a more noticeable effect under a geometric rate.
@@
Extending the proposed approach to multivariate data is an important direction. A natural extension involves ranking multivariate observations based on their norms. Specifically, one could define a suitable norm depending on the task (e.g., Euclidean), compute the norm for each observation, and then apply our univariate ranking method (GoRank) to these one-dimensional values. Observations with the largest norms can then be treated as potential outliers. Following this, a multivariate version of GoTrim can be defined by discarding the top $k=\lfloor\alpha n\rfloor$ observations with the largest norms, and computing the mean of the remaining points using the standard gossip algorithm. Alternatively, one could explore data depths as a generalization of ranking in multivariate settings (see Mosler [2013]). Data depth provides a measure of centrality for multivariate data. While this approach is promising, it would require a more in-depth investigation beyond the scope of the current work, since depth computations usually require to solve computationally demanding optimization problems, just like alternative methods recently designed to define multivariate ranks (e.g. based on optimal transport). Finally, we could simply compute the coordinate-wise trimmed mean by applying GoTrim to each coordinate individually [Pillutla et al., 2022, Yin et al., 2018].
@@
The proposed methods offer a promising foundation for robust decentralized optimization. In particular, GoTrim could be integrated into existing mean-based optimization algorithms to enhance robustness against outliers [Koloskova et al., 2019, Colin et al., 2016, Wu et al., 2018]. Realizing this integration, however, will require further algorithmic and theoretical development. Additionally, our ranking algorithms may prove valuable in extreme value theory, where identifying rare or extreme observations is critical—especially in high-stakes domains such as finance, insurance, and environmental science [Jalalzai et al., 2018].

### v8

@@ Appendix E Robustness Analysis: Breakdown Point of the Partial Trimmed Mean
$\frac{1}{n}\max\left(\left\lfloor m+\frac{1}{2}-\frac{K(\delta)}{\sqrt{t-T}}\right\rfloor,0\right)\leq\varepsilon^{*}_{i}(t)\leq\frac{m}{n}\kern 5.0pt,$
@@
Following the proof in [boyd2006randomized], by repeatedly conditioning, we obtain
@@ Appendix I Large-scale Experiments
Designing faster gossip algorithms is an interesting direction, and we have also been exploring this aspect. One promising approach involves optimizing the edge sampling strategy based on the graph connectivity constant $c$ (see [xiao2004fast, boyd2004fastest]). While this strategy improves gossip performance in standard averaging tasks, our empirical experiments suggest that its impact on the GoRank algorithm is more limited. This is likely due to the relatively small variation in $c$, which is insufficient to significantly affect a rate in 1/ct1/ct—though it has a more noticeable effect under a geometric rate.
@@
Extending the proposed approach to multivariate data is an important direction. A natural extension involves ranking multivariate observations based on their norms. Specifically, one could define a suitable norm depending on the task (e.g., Euclidean), compute the norm for each observation, and then apply our univariate ranking method (GoRank) to these one-dimensional values. Observations with the largest norms can then be treated as potential outliers. Following this, a multivariate version of GoTrim can be defined by discarding the top $k=\lfloor\alpha n\rfloor$ observations with the largest norms, and computing the mean of the remaining points using the standard gossip algorithm. Alternatively, one could explore data depths as a generalization of ranking in multivariate settings (see [mosler2013depth]). Data depth provides a measure of centrality for multivariate data. While this approach is promising, it would require a more in-depth investigation beyond the scope of the current work, since depth computations usually require to solve computationally demanding optimization problems, just like alternative methods recently designed to define multivariate ranks (e.g. based on optimal transport). Finally, we could simply compute the coordinate-wise trimmed mean by applying GoTrim to each coordinate individually [pillutla2022robust, yin2018byzantine].
@@
The proposed methods offer a promising foundation for robust decentralized optimization. In particular, GoTrim could be integrated into existing mean-based optimization algorithms to enhance robustness against outliers [koloskova2019decentralized, colin2016gossip, wu2018review]. Realizing this integration, however, will require further algorithmic and theoretical development. Additionally, our ranking algorithms may prove valuable in extreme value theory, where identifying rare or extreme observations is critical—especially in high-stakes domains such as finance, insurance, and environmental science [jalalzai2018binary].